A coil of annealed copper wire has 820 turns, the average length of which is 9 in. If the diameter of the wire is 32 mils, calculate the total resistance of the coil at 20 degree celsius.
1
Expert's answer
2020-09-07T09:02:57-0400

"Resistance=N\space \rho\frac{l}{A}"


"Area\space of \space cross section=\frac{\pi d^{2}}{4}"


Therefore resistance ="4N\rho\frac{l}{\pi d^{2}}"


Resistivity of annealed copper at "20\degree"is "1.72\times10^{-8}"

length ="9\space inches=0.2286m"

"N=820 \space turns"

Diameter="32\space mils=0.0008128m"


Resistance="4(820)(1.72\times10^{-8})\frac{0.2286}{\pi \times 0.0008128^{2}}=6.2132"


Resistance= 6.2132 "\Omega"